# Distribution Grids May Be a Barrier to Residential Electrification

### v3

@@ DISCUSSION
This paper used bottom-up modeling to estimate the physical and economic impacts of electrification of all residential appliances and personal vehicles on distribution grids in each county of the lower 48 United States. The modeling tools are open-source; the cleaned input data and simulation results are free and public. This paper found that distribution grid reinforcement requirements depend mainly on space heating demand, which can increase peak aggregate demand by a factor of four or more in the coldest counties. Without demand-side management, distribution grid reinforcement could cost \$350 to \$790 billion nationwide, or \$2,800 to \$6,400 per household, but smart electrification could eliminate up to three-quarters of these costs.
@@
This section discusses the sensitivity of nationwide distribution grid reinforcement requirements to several modeling assumptions, including heat pump efficiencies, whether households lower their heating temperature setpoints overnight, heat pump sizing protocols, the choice of weather data, the assumed distribution grid headroom, and the electrification adoption rate. Fig. 7 summarizes the results. The Supplementary Information document discusses the sensitivity analyses in more detail.
Heat pump efficiencies significantly influence grid reinforcement requirements. In cold weather, heat pumps become less efficient and may require electric resistance backup to meet heat demand. The central cases presented above assumed that heat pumps in cold regions met the United States Department of Energy’s Cold-Climate Heat Pump Challenge specifications 38. However, in a sensitivity study where heat pumps in cold regions have efficiencies that are similar to today’s equipment, nationwide distribution grid reinforcement costs increase by 25%.
@@
The sensitivities of grid reinforcement requirements to modeling choices on heat pump efficiency (25%), overnight temperature setpoint adjustments (31%), and heat pump sizing (21%) are more than additive. In a joint sensitivity study including all three of the alternative modeling choices discussed above, nationwide grid reinforcement costs by 86%.
The choice of weather data for the all-electric simulations has relatively little influence on grid reinforcement requirements. In the central cases presented above, we used historical weather data from 2021. An alternative approach would be to use hypothetical future weather data that incorporates the effects of climate change. Climate scientists affiliated with the Intergovernmental Panel on Climate Change have developed frameworks known as the Shared Socioeconomic Pathways (SSP) and Representative Concentration Pathways (RCP) 48. These frameworks provide both quantitative and qualitative descriptions of various societal and environmental development pathways through 2100. For example, in the middle-of-the-road scenario SSP2-RCP4.549, global mean surface temperature rises about 2 C by 2050. In a sensitivity study with 2050 SSP2-RCP4.5 weather data, nationwide grid reinforcement costs decrease by 6%. While climate change increases electricity demand peaks from air conditioning in hot regions, climate change also decreases peaks from space heating in cold regions. In the sensitivity study, the latter effect outweighs the former.
@@
The adoption rate of electric vehicles and heat pumps also influences grid reinforcement requirements. In the central case presented above, we estimated that 100% adoption of electric vehicles and heat pumps would cost \$569 billion nationwide. We ran sensitivity studies, varying the adoption rate from 0% to 100%. Nationwide grid reinforcement costs less than \$1 billion at 10% adoption and a modest \$40 billion at 30% adoption. Above 40% adoption, grid reinforcement requirements increase approximately linearly.
@@
Table 1 summarizes the data sources drawn used in this paper. Distributions of building and appliance properties follow the ResStock database 50. Heat pump coefficient of performance curves come from manufacturer data aggregated by the Northeast Energy Efficiency Partnerships 51. We generate domestic hot water time series from 36. The number of vehicles per household in each county comes from census data 52. We classify personal vehicles as either small (e.g., coupes and sedans) or large (e.g., pickup trucks and sport utility vehicles) and tune the distribution of small and large vehicles in each county to survey data from 53. We tune the driving efficiencies of small and large electric vehicles to the Tesla Model S and Ford F-150 lightning, respectively. For both vehicle types, driving efficiencies vary with outdoor temperatures according to 54; this variation reflects energy use for cabin heating and cooling, as well as temperature-dependent battery chemistry. Commute distances follow census data for each county 55. The Supplementary Information document contains further details on modeling, simulation, and data sources.
@@
Building types and floor areas NREL’s ResStock 25, 56
Business-as-usual appliances NREL’s ResStock 25, 56
Building insulation and air sealing ASHRAE code 57
Design heating and cooling temperatures EnergyStar 58
Heat pump coefficients of performance Northeast Energy Efficiency Partnerships 51
Domestic hot water profiles Building America analysis spreadsheets 36
Commute times Census survey 59
Vehicles per household Census survey 59
Electric vehicle parameters Manufacturer specification and journals 54, 60
Weather data Oikolab 34
Climate zones ASHRAE/IECC 61
Grid upgrade prices Journals 41
@@
where $\eta_{1}$ is the heat pump’s coefficient of performance. Experimental validation of the space heating model can be found in 33. Space cooling works similarly but without resistance backup.
@@
$\underbrace{\frac{\pi_{\text{recurring}}G}{n}\sum_{k=1}^{n}\frac{k(1+i)^{k}}{(1+r)^{k}}}_{\text{Present value of recurring costs}}+\underbrace{\frac{\pi_{\text{capital}}G}{n}\sum_{k=1}^{n}\frac{(1+i)^{k}}{(1+r)^{k}}}_{\text{Present value of capital costs}}.$ (9)
@@
We estimate the grid reinforcement cost using the prices $\pi_{\text{recurring}}$ and $\pi_{\text{capital}}$ from 41. Since 41 does not model the uncertainty associated with these prices, we assume they follow Gaussian distributions with standard deviations of 20% of the mean prices. As the total cost in Eq. (9) is a linear combination of $\pi_{\text{recurring}}$ and $\pi_{\text{capital}}$, the total cost is also Gaussian. We calculate its mean and variance analytically using standard formulas.
@@
The data and code that support this study are available at https://github.com/priyada7/EDGIE.
@@ ACKNOWLEDGMENTS
This work was supported through the Center for High Performance Buildings at Purdue University. The authors thank James Braun, Roy Crawford, Andrew Hjortland, Scott Munns, Wayne Craft, Ronald Domitrovic, Elias Pergantis, and Sophia Evers for helpful discussion.
@@
Document S1. Supplemental experimental procedures, Figures S1–S13.

### v4

@@ DISCUSSION
This paper used bottom-up modeling to estimate the physical and economic impacts of electrification of all residential appliances and personal vehicles on distribution grids in each county of the lower 48 United States. The modeling tools are open-source; the cleaned input data and simulation results are free and public. This paper found that distribution grid reinforcement requirements depend mainly on space heating demand, which can increase peak aggregate demand by a factor of four or more in the coldest counties. Without demand-side management, distribution grid reinforcement could cost \$350 to \$790 billion nationwide, or \$2,800 to \$6,400 per household, but smart electrification could eliminate up to two-thirds of these costs.
@@
This section discusses the sensitivity of nationwide distribution grid reinforcement requirements to several modeling assumptions, including heat pump efficiencies, whether households lower their heating temperature setpoints overnight, heat pump sizing protocols, the choice of weather data, the assumed distribution grid headroom, and the electrification adoption rate. Fig. 7 summarizes the results. The Supplemental Notes discuss the sensitivity analyses in more detail.
Heat pump efficiencies significantly influence grid reinforcement requirements. In cold weather, heat pumps become less efficient and may require electric resistance backup to meet heat demand. The central cases presented above assumed that heat pumps in cold regions met the United States Department of Energy’s Cold-Climate Heat Pump Challenge specifications 37. However, in a sensitivity study where heat pumps in cold regions have efficiencies that are similar to today’s equipment, nationwide distribution grid reinforcement costs increase by 25%.
@@
The sensitivities of grid reinforcement requirements to modeling choices on heat pump efficiency (25%), overnight temperature setpoint adjustments (31%), and heat pump sizing (21%) are more than additive. In a joint sensitivity study including all three of the alternative modeling choices discussed above, nationwide grid reinforcement costs by 87%.
The choice of weather data for the all-electric simulations has relatively little influence on grid reinforcement requirements. In the central cases presented above, we used historical weather data from 2021. An alternative approach would be to use hypothetical future weather data that incorporates the effects of climate change. Climate scientists affiliated with the Intergovernmental Panel on Climate Change have developed frameworks known as the Shared Socioeconomic Pathways (SSP) and Representative Concentration Pathways (RCP) 49. These frameworks provide both quantitative and qualitative descriptions of various societal and environmental development pathways through 2100. For example, in the middle-of-the-road scenario SSP2-RCP4.550, global mean surface temperature rises about 2 C by 2050. In a sensitivity study with 2050 SSP2-RCP4.5 weather data, nationwide grid reinforcement costs decrease by 8%. While climate change increases electricity demand peaks from air conditioning in hot regions, climate change also decreases peaks from space heating in cold regions. In the sensitivity study, the latter effect outweighs the former.
@@
The adoption rate of electric vehicles and heat pumps also influences grid reinforcement requirements. In the central case presented above, we estimated that 100% adoption of electric vehicles and heat pumps would cost \$570 billion nationwide. We ran sensitivity studies, varying the adoption rate from 0% to 100%. Nationwide grid reinforcement costs less than \$1 billion at 10% adoption and a modest \$40 billion at 30% adoption. Above 40% adoption, grid reinforcement requirements increase approximately linearly.
@@ METHODS
Table 1 summarizes the data sources drawn used in this paper. Distributions of building and appliance properties follow the ResStock database 51. Heat pump coefficient of performance curves come from manufacturer data aggregated by the Northeast Energy Efficiency Partnerships 52. We generate domestic hot water time series from 35. The number of vehicles per household in each county comes from census data 53. We classify personal vehicles as either small (e.g., coupes and sedans) or large (e.g., pickup trucks and sport utility vehicles) and tune the distribution of small and large vehicles in each county to survey data from 54. We tune the driving efficiencies of small and large electric vehicles to the Tesla Model S and Ford F-150 lightning, respectively. For both vehicle types, driving efficiencies vary with outdoor temperatures according to 55; this variation reflects energy use for cabin heating and cooling, as well as temperature-dependent battery chemistry. Commute distances follow census data for each county 56. The Supplemental Notes contain further details on modeling, simulation, and data sources.
@@
Building types and floor areas NREL’s ResStock 22, 57
Business-as-usual appliances NREL’s ResStock 22, 57
Building insulation and air sealing ASHRAE code 58
Design heating and cooling temperatures EnergyStar 59
Heat pump coefficients of performance Northeast Energy Efficiency Partnerships 52
Domestic hot water profiles Building America analysis spreadsheets 35
Commute times Census survey 60
Vehicles per household Census survey 60
Electric vehicle parameters Manufacturer specification and journals 55, 61
Weather data Oikolab 33
Climate zones ASHRAE/IECC 62
Grid upgrade prices Journals 40
@@
where $\eta_{1}$ is the heat pump’s coefficient of performance. Experimental validation of the space heating model can be found in 32. Space cooling works similarly but without resistance backup.
@@
[…]
@@
We estimate the grid reinforcement cost using the prices $\pi_{\text{recurring}}$ and $\pi_{\text{capital}}$ from 40. Since 40 does not model the uncertainty associated with these prices, we assume they follow Gaussian distributions with standard deviations of 20% of the mean prices. As the total cost in Eq. (9) is a linear combination of $\pi_{\text{recurring}}$ and $\pi_{\text{capital}}$, the total cost is also Gaussian. We calculate its mean and variance analytically using standard formulas.
@@
The data and code that support this study are available at https://github.com/priyada7/EDGIE and at https://zenodo.org/records/15857150.
@@ ACKNOWLEDGMENTS
This work was supported through the Purdue University Center for High Performance Buildings grant #60. The authors thank James Braun, Roy Crawford, Andrew Hjortland, Scott Munns, Wayne Craft, Ronald Domitrovic, Elias Pergantis, and Sophia Evers for helpful discussion. The authors thank the anonymous peer reviewers for improving the manuscript through their detailed and insightful feedback.
@@
Document S1. Supplemental experimental procedures, Figures S1–S13, Table S1.